# Refining Skewed Perceptions in Vision-Language Models through Visual Representations

### v2

@@ 6 Visual Representations Can Refine the Skewed Perception in VLMs
(see proof in supplementary materials). is then processed through the ERM-trained linear probe to produce the final classification result. For simplicity and ease of implementation, we assume Euclidean orthogonality, as using non-Euclidean orthogonality—specifically, employing the inverse of the covariance matrix in the neighborhood space as the metric for inner product calculations—did not yield significant improvement.
The definition of “background” image varies with the dataset. For the Waterbirds dataset, artificially created with images from Places [29] and CUB [23], we defined a range of “background” conditions from least related (random images from Places) to most related (natural environments like lakes and forests) to specific backgrounds used in Waterbirds. Figure 7 shows examples of background images with varying levels of semantic similarity to the Waterbird sample image’s background, which is used in Table 1. In Table 1, we demonstrate that the experiments using semantically more similar “background” images yields higher WGA in Waterbirds. Besides, transitioning to an ERM-trained linear probe enhances WGA further by focusing more sharply on core features, unlike the “contaminated” text representations from CLIP. Both supervised (knowing the corresponding “background” category, denoted by in Table 1) and unsupervised (without knowing the corresponding “background” category, denoted by ) projections were explored, with the supervised setup serving to illustrate the upper limit of this approach, rather than its practical applicability. Increasing the number of “background” vectors generally improves WGA, but with diminishing returns. The VisualDistiller can achieve the WGA of 82.40% without knowing the background image category (20 random images from nature, ViT), only a few points from supervised DFR’s 85.67% performance.
@@
| a random image from Places | 1 | 42.68% | 90.56% | 48.29% | 90.65% | |||||||
| a random image from same class | 1 | 54.83% | 86.95% | 66.82% | 86.41% | |||||||
| a random image from same subclass | 1 | 57.94% | 87.51% | 71.34% | 81.68% | |||||||
@@
| random images from Places | 10 | 71.81% | 96.06% | 63.08% | 94.08% | |||||||
@@
| random images from nature | 20 | 82.40% | 95.03% | 62.77% | 90.15% | |||||||
@@
| random images from same class | 10 | 87.07% | 93.45% | 73.99% | 89.86% | |||||||
@@
| random images from same subclass | 10 | 87.85% | 93.35% | 72.90% | 89.82% | |||||||
@@
Additionally, we applied VisualDistiller to the CelebA dataset classifying if the celebrity’s hair color is blond. Here, we used images of celebrities without hair as “background” vectors source to mitigate non-hair related features. For examples of “background” image in this experiment, see Figure 8. Despite real-world limitations preventing the exact matching of these “background” conditions, using a set of bald celebrity images proved effective. In Table 2, results show significant improvements in WGA with ERM projections, particularly when using gender-matched bald celebrity images. We observed that the WGA on an ERM linear probe escalated from 47.22%/38.89% (no projection on ViT/ResNet) to 83.88%/83.33% (projecting with a corresponding gender bald image on ViT/ResNet). However, projections using irrelevant or opposite-gender images tended to reduce the WGA gains achieved through gender-matching bald images, highlighting the specificity required for effective “background” vector selection. Although using text-based “background” vectors assisted in refining the projection, the inherent biases within text representations limited their effectiveness compared to image-based projections. More experiments on other CelebA attributes can be found in Table 3. Note that the proposal of VisualDistiller purely aim to validate the effectiveness of visual represention over the text representation, hence we do not seek to benchmark with other methods like supervised DFR.
@@
| an opposite gender image | 66.67% | 85.45% | 66.11% | 87.98% | |||||
| a male and female image | 79.37% | 86.21% | 81.11% | 87.43% | |||||
| ERM | a corresponding gender image | 83.88% | 87.60% | 83.33% | 87.76% | ||||
@@
Minority Group Accuracy | ||||
@@
| an opposite gender image | 95.97% | ||||||
@@
| ERM | a corresponding gender image | 94.85% | |||||
@@
| an opposite gender image | 97.97% | ||||||
@@
| ERM | a corresponding gender image | 94.56% | |||||
@@ 7 Conclusions
In this study, we explored the capabilities of a CLIP model in manipulating a linear probe from multiple perspectives. We showed that the text prompt representations are often tainted by contextual features embedded within the training data. In contrast, visual representations demonstrated greater expressiveness, and targeting specific features within these representations proved highly effective for extracting essential information for downstream tasks. Our straightforward, cost-effective, and potent framework VisualDistiller is intended to generate further insights into the crafting of representations in VLMs and provide the community with a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
Limitations and Broader Impacts.
Mitigating spurious correlations in machine learning models is crucial for developing more reliable and trustworthy AI. Regarding privacy and security risks, these are relatively low in our study, as our work builds upon certified VLMs like CLIP. Looking ahead, our future research will focus on addressing challenges within the broader scope of spurious correlations embedded in VLMs. This includes using a non-linear probe to distill task-specific core features and mitigating bias.

### v3

@@ 6 Visual Representations Can Refine the Skewed Perception in VLCMs
(see proof in supplementary materials). is then processed through the ERM-trained linear probe to produce the final classification result. For simplicity and ease of implementation, we assume Euclidean orthogonality, proven effective in [12, 15]. Using non-Euclidean orthogonality — specifically, employing the inverse of the covariance matrix in the neighborhood space as the metric for inner product calculations — did not yield substantial improvement.
The definition of a “background” image varies with the dataset. For the Waterbirds dataset, artificially created with images from Places [33] and CUB [25], we defined a range of “background” conditions from loosely related (retrieved111Here, “retrieved” refers to the process of encoding all images in the Places dataset using the CLIP ViT-L/14 image encoder and then identifying the top 50 images whose embeddings are most similar to the text embedding using K-nearest neighbors. images from Places using “a photo of indoor or outdoor background”) to more strictly related (retrieved images from Places using “a photo of nature background”) to corresponding backgrounds used in Waterbirds. Figure 7 shows examples of background images with varying levels of semantic similarity, which is used in Table 1. In Table 1, we demonstrate that the experiments using semantically more similar “background” images yield higher WGA in Waterbirds. Besides, transitioning to an ERM-trained linear probe enhances WGA further by focusing more sharply on core features, unlike the “contaminated” text representations from CLIP. Both supervised (knowing the corresponding “background” category, denoted by in Table 1) and unsupervised (without knowing the corresponding “background” category, denoted by ) projections were explored, with the supervised setup serving to illustrate the upper limit of this approach, rather than its practical applicability. Increasing the number of “background” vectors generally improves WGA, but with diminishing returns. The VisualDistiller can achieve the WGA of 82.40% without knowing the background image category (20 random images from nature, ViT), only a few points from supervised DFR’s 85.67% performance. We also observe a minor decrease in the average accuracy of our proposed experiments compared to ERM without projection on the original Waterbirds dataset. This is expected, and widely exists in similar framework [27, 28], as ERM prioritizes overall accuracy at the expense of minority group performance — the main optimization subject of the work.
@@
| an image retrieved by “a photo of indoor or outdoor background” | 1 | 42.68% | 90.56% | 48.29% | 90.65% | |||||||
| an image retrieved by corresponding class text | 1 | 54.83% | 86.95% | 66.82% | 86.41% | |||||||
| an image retrieved by corresponding subclass text | 1 | 57.94% | 87.51% | 71.34% | 81.68% | |||||||
@@
| images retrieved by “a photo of indoor or outdoor background” | 10 | 71.81% | 96.06% | 63.08% | 94.08% | |||||||
@@
| images retrieved by “a photo of nature background” | 20 | 82.40% | 95.03% | 62.77% | 90.15% | |||||||
@@
| images retrieved by corresponding class text | 10 | 87.07% | 93.45% | 73.99% | 89.86% | |||||||
@@
| images retrieved by corresponding subclass text | 10 | 87.85% | 93.35% | 72.90% | 89.82% | |||||||
@@
Additionally, we applied VisualDistiller to the CelebA dataset classifying if the celebrity’s hair color is blond. Here, we used retrieved222Here, “retrieved” refers to the process of encoding all images in the CelebA dataset using the CLIP ViT-L/14 image encoder and then identifying the top 50 images whose embeddings are most similar to the query embedding using K-nearest neighbors. images of celebrities without hair as “background” vector source to mitigate non-hair related features. For examples of “background” images in this experiment, see Figure 8. Despite real-world limitations preventing the exact matching of these “background” conditions, using a set of bald celebrity images proved effective. In Table 2, results show significant improvements in WGA with ERM projections, particularly when using gender-matched bald celebrity images. We observed that the WGA on an ERM linear probe escalated from 47.22%/38.89% (no projection on ViT/ResNet) to 83.88%/83.33% (projecting with a corresponding gender bald image on ViT/ResNet). However, projections using irrelevant or opposite-gender images tended to reduce the WGA gains achieved through gender-matching bald images, highlighting the specificity required for effective “background” vector selection. Although using text-based “background” vectors assisted in refining the projection, the inherent biases within text representations limited their effectiveness compared to image-based projections. More experiments on other CelebA attributes can be found in Table 3. Note that the proposal of VisualDistiller purely aims to validate the effectiveness of visual representation over the text representation, hence we do not seek to benchmark with other methods like supervised DFR.
@@
| an image retrieved by opposite gender text | 66.67% | 85.45% | 66.11% | 87.98% | |||||
| two images retrieved by corresponding and opposite gender text | 79.37% | 86.21% | 81.11% | 87.43% | |||||
| ERM | an image retrieved by corresponding gender text | 83.88% | 87.60% | 83.33% | 87.76% | ||||
@@
Worst Group Accuracy | ||||
@@
| an image retrieved by opposite gender text | 95.97% | ||||||
@@
| ERM | an image retrieved by corresponding gender text | 94.85% | |||||
@@
| an image retrieved by opposite gender text | 97.97% | ||||||
@@
| ERM | an image retrieved by corresponding gender text | 94.56% | |||||
@@ 7 Conclusions
In this study, we explored the capabilities of a CLIP model in manipulating a linear probe from multiple perspectives. We showed that the text prompt representations are often tainted by contextual features embedded within the training data. In contrast, visual representations demonstrated greater expressiveness, and targeting specific features within these representations proved highly effective for extracting essential information for downstream tasks. Our straightforward, cost-effective, and potent framework VisualDistiller is intended to generate further insights into the crafting of representations in VLCMs and provide the community with a more comprehensive understanding of the distinct capabilities and limitations of CLIP’s visual and textual representations.
Broader Impacts.
Mitigating spurious correlations in machine learning models is essential for building reliable and trustworthy AI. For instance, recognizing that text embeddings contain spurious biases helps us approach their use in image retrieval with caution. In future work, we aim to enhance fairness in image retrieval by leveraging distilled image embeddings as a surrogate for query embeddings, reducing bias in retrieved results.